# Optical excitation of bulk plasmons in n-doped InAsSb thin films : investigating the second viscosity in electron gas

### v1

@@ III Improved parameter retrieval
Figure 12: Experimental reflectance data (solid lines) for several thicknesses and the result of a global fit (dashed lines) on all the curves at once for sample #1. The curves are shifted vertically to be more clearly visible. The values provided by the optimization are $n_{0}=5.31\times 10^{18}$ cm3, $m^{*}=0.057\,m_{0}$ and $\xi=2.11\times 10^{-10}$ Pa.s.
IV Discussion
IV.1 The slab as a two mode cavity
The model proposed by Ferrel, which describes the electron gas as somewhat rigid, is appealing due to its simplicity. The picture of a cavity for longitudinal waves is however much more accurate and powerful, and we will detail all its aspects here.
In that framework, it is easy to estimate the position of the resonances. Given the large value of the wavevector for plasmons, even when its component along the interface $k_{x}$ increases, the normal $k_{z}$ component remains essentially unchanged. Whatever the incidence angle, the plasmon can thus be considered to propagate almost perfectly perpendicularly to the interfaces. The resonance condition for such a cavity can be expressed as:
$k=\frac{\ell\pi}{h}$ (9)
where $\ell$ is an non zero integer, $h$ the thickness of the slab and $k$ the plasmon wavevector. Using the lossless dispersion relation for the plasmon, this provides an approximation for the resonance frequencies:
$w_{m}=\sqrt{w_{0}^{2}+\left(\frac{\ell\beta\pi}{h}\right)^{2}}.$ (10)
Despite its simplicity and the fact that the losses are neglected, this formula is quite accurate, as shown in Fig. 13, where the lossless dispersion relation is used to compute the position of the resonance on the spectrum. The spectrum shown is computed with $\xi=0$ to better see the position of the resonances. The figure shows clearly that only the resonance corresponding to an odd value of $\ell$ can be seen. We discuss below how the cavity picture is able to explain also this characteristics.
The accuracy of formula (10) justifies the assumptions that the incidence angle or the losses have little impact on the position of the resonances – except for the fundamental one because the dispersion relation for the plasmon including losses significantly differs for low wavevectors, as shown in Fig. 1.
Of course, the amplitude of each resonance is dependent on the incidence angle. In the Ferrel picture and in the Drude model framework, it cannot be excited in normal incidence because it is a vertical oscillation. In the cavity picture, it is not excited in normal incidence because the electric field associated to the plasmon needs a component perpendicular to the interface to be excited.
The cavity picture allows to understand why the first resonance does not occur exactly at $\omega_{0}$, if $\omega_{0}$ is defined as the frequency for which the real part of the permittivity vanishes. As shown in Figure 13, there is a shift between the position of the first resonance predicted by the hydrodynamic model and by a Drude model, which assumes a rigid electron gas. While this shift is small for such a thick structure it becomes large for thinner structures as shown in Fig. 4 or 6 for instance. Such a large shift is thus a clear sign of a nonlocal response, dictated by the dispersion relation of plasmons. It gives a clear indication on the value of $\beta$. This simple model also allows to understand that, when $\beta$ tends to zero, all the resonances coalesce into a single resonance, corresponding to the Ferrel resonance within the framework of the Drude model.
Figure 13: Top: Simulated reflectance spectrum for a 70 nm thick slab without any imaginary part for $\eta^{2}$ (blue curve) and prediction using the Drude model only (orange curve). Only the odd resonances (grey dashed lines) are visible, while even resonances are absent. A simple cavity formula provides an accurate way to compute the position of these resonances, allowing to better understand all the characteristics of plasmon resonances and how they behave when changing $\beta$, the incidence angle of the thickness of the semiconductor. Bottom: Dispersion curve for plasmons, and cavity resonance conditions on $k$ allowing to compute the resonance frequencies.
The fact that the even resonances of a cavity cannot be excited is a phenomenon that typically arises when a cavity is excited in phase and with the same amplitude from both sides[26], a phenomenon sometimes referred to as coherent perfect absorption[27].
At first glance, the physical situation here does not seem to fit this description because the structure is illuminated from above only. However, the slab functions as a two-mode cavity[28]. The transverse wave also traverses the semiconductor slab, and since we are very close to $\omega_{0}$, the effective refractive index approaches zero, resulting in a vanishing wavevector and a diverging wavelength. To be precise, at the considered incidence angle, the transverse wave is usually evanescent in the medium but has a very large penetration depth. In both scenarios, the slab thickness is small compared to either the wavelength or the penetration depth, allowing it to reach the bottom interface while maintaining the same phase and amplitude. The plasmon that is generated at the bottom interface is then expected to have the same phase as the plasmon generated at the top interface and a comparable amplitude, thus canceling the even resonances of the cavity. We underline that this also leads to the enhancement of odd resonances. This is fundamental in the so called coherent perfect absorption and explains the large absorption cross-section of plasmonic nanocavities[29] because is enhances absorption at resonance by a factor of typically four. Without such a phenomenon, it is likely only the main resonance could have been observed here.
IV.2 Second viscosity of the electron gas
Given the narrow range of frequencies around the screened plasma frequency for which plasmon resonances can be observed, and the uncertainties regarding the imaginary part of $\beta$, it is not possible to determine which theory proposed in earlier works [12, 9, 8] is accurate by monitoring the dependency of the hydrodynamic parameter on frequency. As shown above, a constant imaginary part allows very accurate fits of the experimental data. Moreover, we believe that a physically satisfactory explanation for the additional losses experienced by plasmons, beyond friction with the lattice, is still lacking.
The GNOR model introduces an additional diffusion current into the charge conservation equation. While this has the advantage of producing a supplementary term that can be accounted for by using a dispersive $\eta^{2}$ as above, the assumption of a supplementary diffusion current is a strong, unprecedented, one. It is not widely accepted.
Another approach has thus been proposed, probably inspired by the fact that in conductive 2D systems, electron gas seem to exhibit a viscous behavior[30, 31], well described by a hydrodynamic model. It consists in adding a viscous term to the hydrodynamic model[13], with a shear viscosity term[8]. When this term is taken into account however, the equations (4) are deeply modified. An imaginary part to $\beta$ should ideally correspond to a term proportional to (.$\nabla(\nabla.\mathbf{P_{f}})$ whereas the viscosity adds also a $\Delta\mathbf{P_{f}}$ that potentially lifts the degeneracy of the plasmon leading to two distinct forms, as is sometimes incorrectly assumed[32].
We think however that there is a lot of merit in such an approach. The fact that adding a simple imaginary part to $\beta^{2}$ seems so accurate, as shown above, suggests strongly that a term (.$\nabla(\nabla.\mathbf{P_{f}})$ is missing. A lot of attention has been paid to the shear viscosity, but we would like to underline that another viscosity has been overlooked: the second (or bulk) viscosity of the electron gas[33]. This form of viscosity intervenes when the fluid expands or contracts significantly. It is typically neglected when deriving the Navier-Stokes equations because most fluid flows are essentially incompressible, even if the fluid itself is very compressible, leaving shear as the only source of friction[34]. This assumption, called the Stokes hypothesis, is so common and accurate that the bulk viscosity of a gas is challenging to estimate. Furthermore, monoatomic gases are expected to have a vanishing bulk viscosity. However, bulk viscosity is the dominant source of attenuation for sound in gases, and in such cases, its value has been found to be orders of magnitude larger than that of shear viscosity[35, 36].
In the present case, the shear within the electron gas, where longitudinal waves propagate, is minimal, while the gas is primarily subjected to high-frequency compressions and expansions. This is typically a situation in which the second viscosity can be expected to dominate.
In order to see how the second viscosity can be taken into account, we consider that the electron gas is a newtonian fluid obeying to the following equation[35, 37, 33, 34] :
where $n$ and $m^{*}$ are the electron density and electron effective mass, respectively, $\mathbf{v}$ the electron velocity, $p$ the pressure inside the electron gas, $\mu$ the (shear) viscosity, $\gamma$ a damping factor and $\xi$ the second (or bulk) viscosity.
If, as explained above, shearing is minimal, we can assume that the first viscosity can be neglected. As is classically done to obtain a linear response model, $\mathbf{v}$ is considered a first order term, so that the inertial term $\mathbf{v}\cdot\nabla\mathbf{v}$ can be neglected also. The electron gas density $n$ can be decomposed into a constant term $n_{0}$ and a first-order fluctuation $n_{1}$ so that finally we can write $-n_{0}e\partial_{t}\mathbf{v}\simeq\partial_{t}\,\mathbf{j}$ at the first order. This can also be done for the spatial derivatives of the bulk viscosity term. Taking these steps into account leads to
$-\frac{m^{*}}{e}\,\partial_{t}\mathbf{j}=-\nabla p+\frac{m^{*}}{e}\,\gamma\,\mathbf{j}-n_{0}\,e\,\mathbf{E}-\frac{\xi}{n_{0}\,e}\,\nabla(\nabla\cdot\mathbf{j})$ (12)
The pressure term is given by quantum theories and is a power of $n$ (usually $n^{3}$ is retained[12]). Whatever the exponent, the gradient of the pressure is simply proportional to the gradient of the electron gas density $n$. The nonlocal parameter $\beta$ is introduced at this stage so that
$-\nabla p=-m^{*}\beta^{2}\,\nabla n.$ (13)
The continuity equation can be written
$\nabla\mathbf{j}=-\partial_{t}(-e\,n)=e\,\partial_{t}n_{1}.$ (14)
Since $\mathbf{j}=\partial_{t}\mathbf{P_{f}}$, this equation can be integrated to yield $\nabla\mathbf{P_{f}}=en_{1}$. The pressure term can be finally written
$-\nabla p=-\frac{m^{*}}{e}\beta^{2}\,\nabla(\nabla.\mathbf{P_{f}}).$ (15)
For the second viscosity term, we have:
$-\frac{\xi}{n_{0}e}\,\nabla(\nabla\cdot\mathbf{j})=-\frac{\xi}{n_{0}e}\,\partial_{t}\nabla(\nabla\cdot\mathbf{P_{f}}),$ (16)
$-\left(\beta^{2}-\frac{i\omega\xi}{n_{0}m^{*}}\right)\nabla(\nabla\cdot\mathbf{P_{f}})-(\omega^{2}-i\omega\gamma)\mathbf{P_{f}}=\epsilon_{0}\omega_{p}^{2}\mathbf{E}.$ (17)
Thus, the second viscosity term appears exactly as the imaginary part of $\beta^{2}$ in the classical hydrodynamic model, with the form given above.
Formally, this is equivalent to the GNOR model, except that we do not need to assume the existence of additional diffusion currents, which is questionable and has an impact on the real part of $\eta^{2}$ also (the difference is however two orders of magnitude smaller than $\beta^{2}$ in our case). This finally would provide convincing explanation of why the negative imaginary part of $\beta^{2}$ acts as an additional source of dissipation. This is particularly obvious when writing the dispersion relation for the plasmon that now becomes:
The magnitude of the supplementary losses is expected to increase with the wavevector, and thus with the order of the resonance.
Using the improved parameter retrieval method described above, we obtained $\xi$ values of $2.11\times 10^{-10}$ Pa.s for sample #1 with a doping of $5.31\times 10^{18}$ cm-3 and $5.7\times 10^{-10}$ Pa.s for sample #2 at higher frequencies due to a larger doping level of approximately $1.2\times 10^{19}$ cm-3. Despite the difficulty in comparing these values due to differing electron gas densities, they both demonstrate a notable consistency.
These values of the bulk viscosity are particularly low compared to polyatomic gases, for which measurements have been made[35]. Assuming an electron gas shares some characteristics with a monoatomic gas, such a low value is reasonable: monoatomic gases are the only ones that are expected to satisfy Stokes’ hypothesis, leading to zero bulk viscosity[36], because they lack rotational or vibrational freedom degrees. The fact that an electron gas nonetheless exhibits a second viscosity suggests that a microscopic theory for bulk viscosity in an electron gas is notably lacking. We suspect the states available for the electrons in the conduction band play a role similar to the available freedom degrees in polyatomic gases. Additionally, although $\xi$ itself is low, the electron gas in a doped semiconductor is highly tenuous, so the ratio of $\xi$ to mass density remains high. This explains why it plays such an important role here.
V Conclusion and perspectives
@@
We have proposed a setup that maximizes the visibility of plasmon resonances in highly doped semiconductors. Combined with the numerical methods detailed in this work, our findings show that the hydrodynamic model, with a complex nonlocal parameter $\beta^{2}$, is remarkably sufficient to capture the optical response of these materials with a high level of accuracy. This suggests that no additional mechanisms or adjustments are necessary beyond this model to fully describe the observed phenomena. We have also identified a physically consistent origin for the imaginary part of $\beta^{2}$, grounded in the bulk viscosity (or second viscosity) of the electron gas, offering a convincing and straightforward justification for the well-known GNOR model[9, 38, 39].
@@
We have also introduced a retrieval technique based on an adapted cost function (allowing to quantify the distance between the experimental data and the simulations) and the use of a global optimization algorithm. In the theoretical framework we propose, this technique allows for the estimation of the doping level, the effective mass of electrons and even the bulk viscosity of the electron gas with a single optical measurement - without any abacus. Additionally, we have shown that these parameters can be retrieved manually with satisfactory accuracy, provided that the resonances are clearly identified, making the approach accessible even without advanced computational tools. These parameters can be considered sufficient to fully characterize the optical response of n-doped semiconductors.
@@
This is particularly advantageous because many numerical methods have been developed to simulate the behavior of nonlocal materials using the hydrodynamic model, even in complex geometries[40, 41, 42]. While plasmon resonances have a strong fundamental interest, it is also crucial to highlight that spatial dispersion plays an important but distinct role below the epsilon-near-zero (ENZ) frequency. In this frequency range, highly doped semiconductors emerge as exceptional plasmonic materials capable of supporting surface waves or more complex guided modes with high effective indices (slow light), which can be exploited to design sensors for detecting and characterizing even small quantities of molecules[43, 44]. Previous studies have demonstrated that the slower the plasmonic guided mode[45], the more pronounced the impact of spatial dispersion[16, 46, 47, 39]. This effect is particularly significant near the ENZ frequency and in the presence of high refractive index materials[21], as commonly observed in the IR range where semiconductors exhibit large permittivities. Therefore, we anticipate that nonlocal effects will remain significant even below the ENZ frequency, underscoring the importance of having suitable simulation tools and parameter retrieval techniques.
Overall, spatial dispersion is often overlooked, not only in the study of doped semiconductors and metals but also in ENZ materials more generally. In these materials, which are typically treated as purely local, predictions of unusual and surprising effects have been made[48]. However, as discussed here, relying on a local description involves assuming a somewhat rigid behavior of the electron gas around the ENZ frequency, effectively implying an infinite propagation speed for longitudinal waves. The inclusion of spatial dispersion may alter these predictions and lead to more accurate conclusions.
@@
Despite the challenges in observing nonlocal effects in metals within the visible range, significant efforts have been made over the past decade to develop theoretical and numerical tools and to design structures that maximize these effects. It is striking how these advancements are now proving to be highly relevant for highly doped semiconductors. Given the crucial role of spatial dispersion in these materials, as discussed above, it seems likely that spatial dispersion will soon become a standard approach for describing their optical response. We are pleased to see this area of research transition from a purely theoretical curiosity to a practical, widely-adopted tool and are convinced that this shift will help to answer the remaining fundamental questions.

### v2

@@ III Improved parameter retrieval
Figure 13: Experimental reflectance data (solid lines) for several thicknesses and the result of a global fit (dashed lines) on all the curves at once for sample #1. The curves are shifted vertically to be more clearly visible.
Beyond the technical aspects of parameter retrieval, our results raise fundamental questions about the nature of electron transport in semiconductors. The observation of significant bulk viscosity in electron gases is surprising. Stokes’ hypothesis —that bulk viscosity can be neglected— is widely adopted in fluid dynamics, and theoretical studies suggest it should be particularly valid for monoatomic gases lacking internal degrees of freedom[graves1999bulk]. The electron gas, being fundamentally a collection of identical particles, might be expected to behave similarly[ciraci2017current, conti1999elasticity]. We suspect the available states in the conduction band play a role analogous to the internal degrees of freedom in polyatomic gases, though the microscopic theory connecting band structure to bulk viscosity in electron gases remains to be developed.
The GNOR model and shear viscosity approaches lead to incompatible frequency dependencies for the nonlocal response[de2018viscoelastic]. While our results appear to favor the GNOR framework through the identification of a bulk viscosity term, the full picture remains incomplete. Notably, we observe systematic discrepancies in higher-order resonances: their profiles cannot be accurately reproduced with a constant $\xi$. This suggests a frequency-dependent bulk viscosity $\xi(\omega)$ that would decrease with increasing frequency. This suggested decrease of $\xi$ with frequency aligns with the typical behavior of bulk viscosity in complex fluids, where high-frequency oscillations progressively decouple from slower relaxation processes. This points toward the need for models that go beyond simple constant coefficients, though more controlled experiments — particularly with better-defined incidence angles and precise sample thickness measurements — will be essential to fully characterize this dispersion[ciraci2017current] and advance our theoretical understanding of nonlocal effects in electron gases.
IV Conclusion and perspectives
@@
We have proposed a setup that maximizes the visibility of plasmon resonances in highly doped semiconductors. Combined with the numerical methods detailed in this work, our findings show that the hydrodynamic model, with a complex nonlocal parameter $\beta^{2}$, is remarkably sufficient to capture the optical response of these materials with a high level of accuracy. We have also identified a physically consistent origin for the imaginary part of $\beta^{2}$, grounded in the bulk viscosity (or second viscosity) of the electron gas, offering a convincing and straightforward justification for the well-known GNOR model[mortensen2014generalized, Raza2015, boroviks2022extremely].
@@
We have also introduced a retrieval technique based on an adapted cost function and the use of a global optimization algorithm. In the theoretical framework we propose, this technique could allow for the estimation of the doping level, the effective mass of electrons, the repulsion between electrons and even the bulk viscosity of the electron gas with a single optical measurement - without any abacus. Additionally, we have shown that these parameters can be retrieved manually with satisfactory accuracy, provided that the resonances are clearly identified, making the approach accessible even without advanced computational tools.
@@
While plasmon resonances have a strong fundamental interest, it is also crucial to highlight that spatial dispersion plays an important but distinct role below the epsilon-near-zero (ENZ) frequency. In this frequency range, highly doped semiconductors emerge as exceptional plasmonic materials capable of supporting surface waves or more complex guided modes with high effective indices (slow light), which can be exploited to design sensors for detecting and characterizing even small quantities of molecules[smaali2021reshaping, paggi2023over]. Previous studies have demonstrated that the slower the plasmonic guided mode[ajib2019energy], the more pronounced the impact of spatial dispersion[Moreau2013, raza2013nonlocal, khalid2021influence, boroviks2022extremely]. This effect is particularly significant near the ENZ frequency and in the presence of high refractive index materials[pitelet2019influence], as commonly observed in the IR range where semiconductors exhibit large permittivities. Therefore, we anticipate that nonlocal effects will remain significant even below the ENZ frequency, underscoring the importance of having suitable simulation tools and parameter retrieval techniques.
@@
Despite the challenges in observing nonlocal effects in metals within the visible range, significant efforts have been made over the past decade to develop theoretical and numerical tools[Toscano2012, schmitt2016dgtd, kwiecien2023nonlocal] and to design structures that maximize these effects. It is striking how these advancements are now proving to be highly relevant for highly doped semiconductors. Given the crucial role of spatial dispersion in these materials, as discussed above, it seems likely that spatial dispersion will soon become a standard approach for describing their optical response. We are pleased to see this area of research transition from a purely theoretical curiosity to a practical, widely-adopted tool and are convinced that this shift will help to answer the remaining fundamental questions.